# Device-Native Autonomous Agents for Privacy-Preserving Negotiations

### v2

@@ I Introduction
Insurance and business-to-business (B2B) commerce typically rely on static pricing models, offering customers fixed quotes that do not allow for negotiation. For example, health insurance applicants cannot trade higher deductibles for lower premiums, and procurement managers cannot negotiate bulk discounts in real time. Although manual bargaining is possible, it remains slow, inconsistent, and generally limited to high-value transactions, which leads to rigid, often suboptimal pricing decisions.
@@ II-B Multi-Agent Systems and Negotiation
Zero-knowledge proofs facilitate private computation without disclosing underlying data [3]. zk-SNARKs [5] enable efficient proof generation suitable for practical deployment. Recent research applies these techniques to machine learning [9]. The present work integrates them into negotiation protocols. This allows agents to demonstrate constraint satisfaction without revealing the constraints themselves.
@@ III-A Agentic Workflow Overview
In the Context Expansion, relevant information is retrieved from both short-term memory (STM) and long-term memory (LTM). STM stores the current negotiation state, whereas LTM contains records of past transactions, learned preferences, and domain-specific knowledge. The Intent Understanding interprets the negotiation goal within this expanded context, identifies the negotiation type, and anticipates counterparty expectations.
@@
Real-Time Monitoring continuously tracks negotiation progress against the planned strategy. When counterparty behavior deviates from expected patterns, the monitoring system initiates replanning. The workflow concludes with Outcome Evaluation, which assesses the final agreement, updates memory with lessons learned, and generates feedback to inform future negotiations.
@@ IV-A Experimental Setup
Scenarios. 3 complexity levels are established for each domain. Insurance scenarios comprise:
(1) simple premium negotiation with single-parameter adjustment,
(2) medium coverage adjustment involving deductible-premium tradeoffs, and
(3) complex claims settlement with multiple disputed items.
B2B scenarios comprise:
(1) simple bulk pricing for single-product orders,
(2) medium contract renewal with volume commitments, and
(3) complex multi-term procurement involving delivery, payment, and warranty terms.
Device Tiers. 3 device configurations are evaluated. High-end devices are equipped with an 8-core CPU, 16GB of RAM, and a dedicated NPU, simulating the performance of flagship smartphones and laptops. Mid-range devices include a 4-core CPU and 8GB of RAM without an NPU, which represents typical consumer hardware. Low-end devices have a 2-core CPU and 4GB of RAM, which are typical of budget devices and older hardware.
@@
Baselines. 3 baseline configurations are compared. Cloud-Only performs all processing on remote servers. Device-Only executes all processing locally without cloud access. Naive Edge employs simple threshold-based offloading without optimization.
@@ IV-B Component Performance
The compression-latency tradeoff reflects computational resources. High-end devices use advanced algorithms while low-end devices employ aggressive pruning.
@@
User studies demonstrate a 27% trust improvement when decision trails are available, with minimal computational overhead.
World Model Distillation. Table IV compares teacher and student models. Model size reduction enables device deployment. The 6% accuracy loss is acceptable given 5.5× latency improvement on high-end devices.
@@ IV-C End-to-End Performance
Figure 3 shows success rate decline with complexity. Insurance exhibits steeper degradation (5 points) than B2B (4 points), reflecting its more contentious nature. Both domains maintain >>83% success.
@@ IV-E Ablation Study
Feasibility Check Removal. Success rate decreases by 8% because agents attempt infeasible negotiations. Latency improves by 12% as failures are detected more quickly, but overall user experience deteriorates.
@@
Offloading Removal. Naive routing increases latency by 24% and privacy leakage by 14%.
Safety Critic Removal. Success rate decreases by 9% because agents accept unfair agreements. Although latency improves by 10%, the benefits of preventing failures outweigh the associated overhead.
@@ VI Limitations
System Constraints. The system currently supports only bilateral negotiations [13, 16]. Multi-party scenarios and adversarial counterparties require additional safeguards. Device limitations restrict the reasoning depth of the 500M-parameter model [6] compared to cloud-scale models.
Behavioral Privacy Leakage. While the system achieves a 94% reduction in cryptographic data exposure, zero-knowledge proofs [3] do not prevent behavioral inference attacks [8]. Adversaries observing negotiation dynamics may infer constraint proximity. Current metrics quantify only cryptographic leakage (Sec. IV). Mitigation strategies, such as differential privacy for timing and randomized concession schedules, are identified as directions for future work.

### v3

@@ I Introduction
Insurance and business-to-business (B2B) commerce typically rely on static pricing models, offering customers fixed quotes that do not allow for negotiation. For example, health insurance applicants cannot trade higher deductibles for lower premiums, and procurement managers cannot negotiate bulk discounts in real time. Although manual bargaining is possible, it remains slow, inconsistent, and limited to high-value transactions, leading to rigid, often suboptimal pricing decisions.
@@ II-B Multi-Agent Systems and Negotiation
Zero-knowledge proofs [3] enable private computation without disclosing data. zk-SNARKs [5] provide efficient proof generation. This work integrates them into negotiation protocols to demonstrate constraint satisfaction.
Privacy attacks on machine learning systems extend beyond cryptographic leakage. Timing attacks exploit computational patterns to infer sensitive information [8]. Model inversion and membership inference attacks can reconstruct private inputs. While zero-knowledge proofs mitigate direct data exposure, behavioral inference remains a challenge (Sec. VI).
@@ III-A Agentic Workflow Overview
Context Expansion retrieves information from short-term memory (STM), such as the current negotiation state, and long-term memory (LTM), including past transactions, preferences, and domain knowledge. Intent Understanding interprets the negotiation goal within this context, identifies the negotiation type, and anticipates the counterparty’s expectations.
@@
Real-Time Monitoring tracks progress against the planned strategy and initiates replanning when counterparty behavior deviates. Outcome Evaluation assesses the final agreement, updates memory with lessons learned, and generates feedback for future negotiations.
@@ IV-A Experimental Setup
Scenarios. 3 complexity levels are established for each domain. Insurance scenarios comprise: (1) simple premium negotiation with single-parameter adjustment, (2) medium coverage adjustment involving deductible-premium tradeoffs, and (3) complex claims settlement with multiple disputed items.
B2B scenarios comprise: (1) simple bulk pricing for single-product orders, (2) medium contract renewal with volume commitments, and (3) complex multi-term procurement involving delivery, payment, and warranty terms.
Device Tiers. 3 device configurations are evaluated. High-end devices are equipped with an 8-core CPU, 16GB RAM, and dedicated NPU. Mid-range devices include a 4-core CPU and 8GB RAM without an NPU. Low-end devices have a 2-core CPU and 4GB RAM.
@@
Baselines. 3 baseline configurations are compared. Cloud-Only performs all processing on remote servers. Device-Only executes all processing locally without cloud access. Naive Edge employs simple, unoptimized threshold-based offloading.
@@ IV-B Component Performance
User studies show a 27% increase in trust when decision trails are available.
World Model Distillation. Table IV compares teacher and student models. Model size reduction enables device deployment. The 6% accuracy loss is acceptable given 5.5× latency improvement on high-end devices. Trade-offs become critical in complex negotiations. Accuracy decreases to 82% in multi-parameter scenarios that require deep reasoning. Simple negotiations retain 91% accuracy. When local confidence falls below 0.7, the system routes complex cases to the cloud, thereby balancing privacy and performance.
@@ IV-C End-to-End Performance
Figure 3 shows success rate decline with complexity. Insurance exhibits steeper degradation (5 points) than B2B (4 points). Both domains maintain >>83% success.
@@ IV-E Ablation Study
Feasibility Check Removal. Success rate drops by 8% because agents attempt infeasible negotiations.
@@
Safety Critic Removal. Success rate decreases by 9% as agents accept unfair agreements.
@@ VI Limitations
System Constraints. The system currently supports only bilateral negotiations [12, 15]. Multi-party scenarios require coordination mechanisms beyond pairwise protocols. Adversarial counterparties may exploit predictable agent behavior through strategic delays or extreme offers. The safety critic (Sec. III-G) assumes rational counterparty behavior and may fail against deliberate manipulation. Device limitations restrict the reasoning depth of the 500M-parameter model [6] compared to cloud-scale models.
Behavioral Privacy Leakage. While the system achieves a 94% reduction in cryptographic data exposure, zero-knowledge proofs [3] do not prevent behavioral inference attacks [8]. Adversaries observing negotiation dynamics may infer constraint proximity through timing patterns and concession sequences. Rapid responses signal acceptable terms. Large initial concessions followed by smaller increments reveal reservation prices. Current metrics quantify only cryptographic leakage (Sec. IV). Mitigation strategies, such as differential privacy for timing and randomized concession schedules, are identified as directions for future work.